A 60 foot by 80 foot rectangular walk in a park surrounds a flower bed. If the walk is of uniform width and its area is equal to the area of the flower bed, how wide is the walk? Set up a quadratic equation, factor, and solve.
Answer by mananth(16949)

width of walk = 60 feet
length = 80 feet
..
Area = 4800 ft^2
..
uniform width w
..
GARDEN dimensions
width = 60-2w
length = 80-2w
...
Area of garden = (60-2w)(80-2w)
FOIL
Area = 4800-120w-160w+4w^2
Area = 4w^2-280w+4800
..
Total area / area of garden = 2
4800 / 4w^2-280w+4800 = 2
4800/2 = 4w^2-280w+4800
2400 =4w^2-280w+4800
4w^2-280w+4800-2400=0
4w^2-280w+2400=0
/4
w^2-70w+600=0
w^2-60w-10w+600=0
w(w-60)-10(w-60)=0
(w-60)(w-10)=0
w= 60 OR 10
ignore 60
..
width = 10 feet
...
m.ananth@hotmail.ca